# A Comprehensive Survey on Continual Learning in Generative Models

### v1

@@ 5.1 Problem Setup
Lifelong learning [29, 238] in vision language action extends the concept of traditional robot action control framework by requiring one robot to continually acquire and adapt across a sequence of tasks $\{T^{1},\dots,T^{K}\}$ over its lifespan while retaining knowledge on previous tasks. This continual learning problem for robots can be formulated as a finite-horizon Markov Decision Process (MDP): $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},H,\mu_{0},R)$, where $\mathcal{S}$ represents the state space, $\mathcal{A}$ is the action space, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the transition function, $H$ is the maximum horizon for each episode of a task, $\mu_{0}$ is the initial state distribution, and $R:\mathcal{S}\times\mathcal{A}\rightarrow R$ is the reward function. During the whole training procedure, each task $T^{k}$ trains the policy $\pi$ from last task characterized by its own initial state distribution $\mu_{0}^{k}$, while $\mathcal{S},\mathcal{A},\mathcal{T}$ and $H$ remain unchanged across all tasks.
Training objectives. Continual learning in vision language action model sequentially trains one single policy $\pi$ through multiple tasks. For each task $T^{k},k\in\{1,\cdots,K\}$, the robot is trained on a dataset of $N$ demonstrations $D^{k}=\{\tau^{k}_{i}\}_{i=1}^{N}$ with a corresponding language description $l^{k}$ guiding the goal of the robot. Each trajectory $\tau^{k}_{i}$ in expert demonstrations with length $L^{k}<H$ consists of state-action pairs $\{(s_{t}^{k},a_{t}^{k})\}_{t=1}^{L^{k}}$ sampled from policy $\pi$ conditioned with task $T^{k}$. The policy is trained with a behavioral cloning loss [239] that aims to mimic the action demonstrations in the dataset:
@@
Benchmarks. To evaluate the performance of robots in continual learning, many benchmarks have been constructed in continual learning scenarios from reinforcement learning to vision language action [240, 241, 119]. Continual World [240] builds a continual reinforcement learning benchmark on robotic manipulation tasks. It provides a realistic, computationally accessible testbed that highlights the limitations of existing continual learning methods in reinforcement learning (RL) settings, which often struggle with forward transfer despite mitigating forgetting. LIBERO [119] builds imitation learning environment for robots to analyze the continual learning problem of robots from many factors, such as algorithms, architectures, task sequences, variations and so on, which provides a comprehensive environment for subsequent work.
@@ 5.2 Architecture-based Approach
Architecture-based approaches explore continual learning on architecture design. One line of work focus on exploiting delicate architectures with skill learning to equip the model with skill distinction, i.e., merging and updating old skills while identifying new skills [242, 121, 120, 122] or construct extensible modular for skill stream [243, 244]. LOTUS [120] designs a hierarchical imitation learning framework that builds an ever-growing skill library through unsupervised discovery. It segments human demonstrations using open-vocabulary vision models to identify reusable skills, incrementally clusters them to update existing skills or add novel ones, and trains a meta-controller to flexibly compose skills. By combining continual skill discovery with experience replay, Lotus outperforms fixed-skill baselines in lifelong manipulation. IsCiL [245] incorporates a framework that incrementally learns shareable skills through a prototype-based memory system to enable adaptation in dynamic environments. It maps demonstrations to state embeddings, retrieves relevant skills via skill prototypes, and employs parameter-efficient adapters to preserve knowledge. LEGION [122] presents a framework for lifelong reinforcement learning for robots that integrates Bayesian non-parametric model clustering with language embeddings to preserve learned skills indefinitely and semantically align task understanding. The method employs a Dirichlet Process Mixture Model (DPMM) to dynamically organize knowledge and leverages language-conditioned representations to facilitate flexible skill composition, allowing continual accumulation and reuse of knowledge across robotic manipulation tasks without performance degradation. RLCM [246] formulates a framework for modular lifelong RL that leverages neural composition. The policies are dynamically constructed by combining learned modules with a two-phase learning process. During online exploration, modules are selected and fine-tuned for new tasks, while offline batch RL with experience replay consolidates knowledge into shared modules to retain past capabilities.
Another line of structure design attempts to enhance robot action with the instruction from large language models like planning [247], skill demonstration [248, 249] and so on. They use pre-trained large language models as a knowledge base and retrieve user requests with textual prompts to provide prior knowledge of skills, thereby acquiring new skills and reusing old tasks for knowledge transfer. Due to data and time efficiency exhibited by trained large language model, they show strengths and promise in many scenarios.
@@ 5.3 Regularization-based Approach
Regularization methods for continual learning usually try to keep important parameter/feature unchanged and impose feature-level or parameter-level regularization during training, therefore maintaining ability for past skills [250, 123, 124, 245]. CAMA [123] introduces Confidence-Aware Moving Average, a method that dynamically updates stored knowledge (logits) using confidence scores to balance past and current information without requiring task boundaries. PolyTask [250] proposes a method combining task-specific policy training with behavior distillation, which consolidates individual policies into a single model using offline data. With demonstration-guided reinforcement learning and distilling policy outputs, it handles continual actions and lifelong learning by avoiding interference between tasks. M2Distill [124] proposes a multi-modal distillation framework that preserves consistency across vision, language, and action modalities between incremental learning steps. It exploits: (1) multi-modal feature distillation to constrain latent space drift; (2) policy distillation to maintain alignment of Gaussian mixture model (GMM) action distributions. The approach ensures skill retention without storing prior data, advancing lifelong learning for real-world robotic applications. ELIRL [251] introduces Lifelong Inverse Reinforcement Learning (Lifelong IRL), addressing the challenge of learning multiple sequential tasks and propose Efficient Lifelong IRL to enable knowledge transfer across tasks, improving performance while minimizing user-provided data.
@@ 5.4 Replay-based Approach
Due to the spatial and temporal complexity of robotic manipulation in various and dynamic scenarios, it has become common practice to store part of previous samples and replay them [252, 37] during training procedure to retain knowledge of previous tasks [125, 122, 243, 126]. iManip [125] addresses the challenge of catastrophic forgetting in robotic manipulation. Given that traditional methods fail to retain previously learned skills, ignoring temporal complexity and action complexity, the introduced framework integrates a temporal replay strategy to preserve task continuity and an extendable PerceiverIO [253] architecture with adaptable action prompts. Experimental results highlight the efficiency, scalability, and robustness across complex, long-horizon tasks. Decision-RWKV [243] exploits a novel sequence modeling framework that integrates the efficiency of the Receptance Weighted Key Value (RWKV) with experience replay mechanisms. It leverages linear complexity and memory retention capabilities of RWKV to enable efficient lifelong learning across diverse manipulation tasks, and the experience replay combines a replay buffer for historical experiences reuse while acquiring new skills. RWLA [126] combines retrieval-based local adaptation with a selective weighting mechanism to restore forgotten skills. The method retrieves relevant past demonstrations from a memory buffer, identifies critical failure-prone segments, and prioritizes these segments during local fine-tuning, which allows robots to adapt continually while maintaining performance across diverse manipulation tasks.
@@ 6.1 Problem Setup
where $z_{t}$ denotes the latent code of an image with the new concept, and $z_{t}^{p}$ corresponds to a class-specific image used for prior preservation. $c$ and $c^{p}$ are the respective conditioning vectors; for instance, $c$ may be "a photo of [$\mathcal{V}^{*}$] dog", while $c^{p}$ is "a photo of dog". The special token [$\mathcal{V}^{*}$] is a unique identifier representing the user-specific concept learned from the provided images. $\lambda$ controls the regularization strength.
Continual learning in diffusion models aims to go beyond single-concept personalization by enabling the incremental acquisition of new concepts, while avoiding catastrophic forgetting and eliminating the need to retrain on previous data. Specifically, as shown in Figure 9, assume a user intends to continuously generate content based on newly introduced specific concepts. The generation tasks are defined as $\mathcal{T}={\{\mathcal{T}^{k}}\}_{k=1}^{K}$ with corresponding datasets ${\{\mathcal{D}^{k}}\}_{k=1}^{K}$. Each $k$-th generation task consists of $n_{k}$ pairs of concept images $x_{i}^{k}$ and prompts $c_{i}^{k}$, where $K$ denotes the total number of tasks, and $n_{k}$ is typically in the range of 3 to 5 in accordance with existing studies. Accordingly, to learn the $k$-th personalized concept, the model is optimized using the following training loss:
@@ 6.3 Regularization-based Approach
MuseumMaker [128] aims to retain knowledge of past learned styles while incorporating new customized styles. To mitigate catastrophic overfitting and maintain the integrity of original concepts in diffusion models, it introduces a style distillation loss that captures disentangled style representations. Furthermore, it proposes a dual regularization scheme for LoRA training, which constrains model optimization from both the weight and feature perspectives to achieve balanced knowledge integration. LFS-Diffusion [129] proposes a data-free knowledge distillation strategy applied across diffusion time steps during training, aiming to constrain the model from overfitting to current personalized concepts and thereby preserve the original generative capabilities. During inference, an in-context generation mechanism is introduced to mitigate forgetting of previously learned knowledge and ensure better retention of past concepts. CIDM [130] introduces a concept consolidation loss that employs learnable layer-wise concept tokens and an orthogonal subspace regularizer to disentangle task-specific and task-shared knowledge. During inference, an elastic weight aggregation module merges low-rank weights of previously learned concepts based on their importance. To alleviate concept neglect in multi-concept composition, CIDM further proposes a context-controllable synthesis strategy that enhances regional feature expressiveness via layer-wise textual embeddings and guides generation through region-aware noise estimation. ConceptGuard [131] introduces a set of lightweight yet effective strategies to mitigate forgetting in continual concept learning for diffusion models. It employs shift embeddings for dynamic concept updates, concept-binding prompts to disentangle and compose concepts over time, and memory preservation regularization to slow down harmful parameter drift. Additionally, a priority queue adaptively manages concept replay based on their importance and temporal order. Jha et al. [132] utilize diffusion classifier scores as implicit class-conditional density estimators. This method simultaneously regularizes both parameter-space and function-space consolidation, enabling effective adaptation to evolving user concepts.
@@ 7 Future Directions
Learning Paradigm for Continual Learning. While conventional continual learning methods predominantly rely on supervised fine-tuning (SFT) for incremental knowledge acquisition, emerging reinforcement learning (RL)-based training paradigms, particularly exemplified by GRPO [272], have demonstrated stronger generalization capabilities across a variety of model architectures. This performance advantage has been consistently observed not only in large language models [273, 274, 275], but also in multimodal large language models [276, 277, 278, 279] and vision-language action models [280, 281], suggesting that RL-based methods may serve as a more robust alternative to traditional SFT strategies in continual learning settings. Recent studies [282] have revealed a fundamental difference between the two paradigms: SFT tends to overfit specific training distributions through pattern memorization, whereas RL emphasizes structural comprehension of task objectives, enabling better generalization to novel scenarios. This distinction provides strong motivation for exploring continual learning within an RL framework, which may naturally support long-term adaptation, task abstraction, and reinforcement-driven retention strategies.
@@
Self-generation for Continual Learning. A core challenge in continual learning lies in the inaccessibility of previous task data. In discriminative models, a widely adopted solution involves training auxiliary generative models (e.g.,, GANs [283] or diffusion models) to synthesize pseudo-samples of prior tasks, thereby mitigating catastrophic forgetting [284, 285, 286, 287]. While effective, this approach introduces significant computational overhead and may suffer from distributional mismatch between generated and original data. In contrast, generative models offer a more integrated solution by leveraging their inherent capacity to reconstruct and regenerate data. These models can autonomously replay past task distributions during the learning of new tasks, without the need for separate generative modules. This capability offers two major benefits: reducing additional architectural complexity and achieving better alignment with the original data distribution. As a result, self-generation in generative models presents a promising direction for building more efficient, scalable, and distributionally consistent continual learning systems.
@@
Continual Learning in Large Scale Models. Most existing studies on continual learning for generative models have focused on small- to medium-scale architectures (e.g., models with up to 7B and 13B parameters). However, emerging evidence [288, 289] suggests that continual learning behaviors can differ substantially as model scale increases. This observation highlights two critical research directions. First, there is a pressing need to establish specialized evaluation benchmarks that mitigate information leakage risks [290], particularly given the large and often proprietary datasets used to train large-scale models. Second, it is essential to explore parameter-efficient knowledge integration strategies—such as Chain-of-Thought reasoning (CoT)[162, 291], In-Context Learning (ICL)[292, 293, 294], Agents [295, 296, 297], and Retrieval-Augmented Generation (RAG) [298, 299]—as conventional fine-tuning becomes increasingly infeasible for models with extensive parameter counts.
@@
Continual Learning for Multimodal Generation Beyond Vision and Language. While most existing research in continual generative learning focuses on vision-language tasks, real-world multimodal systems are increasingly incorporating richer modalities such as audio [300], 3D scenes [301], and motion data [302]. In such settings, modality-continual learning becomes a critical yet underexplored challenge. For example, a pretrained MLLM may be required to incorporate speech input or generate spoken responses. Without appropriate continual learning mechanisms, learning the new audio modality could result in severe forgetting of previously acquired vision-language capabilities. Similarly, in VLA models deployed in robotics or embodied agents, the integration of new sensor modalities (e.g., depth and haptic signals) necessitates stable expansion of generative capacity without overwriting existing knowledge. These scenarios highlight the need for continual learning methods capable of integrating new modalities while preserving alignment and generative quality across previously learned ones.

### v2

@@ 5.1 Problem Setup
Lifelong learning [30, 236] in vision language action extends the concept of the traditional robot action control framework by requiring one robot to continually acquire and adapt across a sequence of tasks $\{T^{1},\dots,T^{K}\}$ over its lifespan while retaining knowledge of previous tasks. This continual learning problem for robots can be formulated as a finite-horizon Markov Decision Process (MDP): $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},H,\mu_{0},R)$, where $\mathcal{S}$ represents the state space, $\mathcal{A}$ is the action space, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the transition function, $H$ is the maximum horizon for each episode of a task, $\mu_{0}$ is the initial state distribution, and $R:\mathcal{S}\times\mathcal{A}\rightarrow R$ is the reward function. During the whole training procedure, each task $T^{k}$ trains the policy $\pi$ from the last task characterized by its own initial state distribution $\mu_{0}^{k}$, while $\mathcal{S},\mathcal{A},\mathcal{T}$ and $H$ remain unchanged across all tasks.
Training objectives. Continual learning in the vision-language action model sequentially trains one single policy $\pi$ through multiple tasks. For each task $T^{k},k\in\{1,\cdots,K\}$, the robot is trained on a dataset of $N$ demonstrations $D^{k}=\{\tau^{k}_{i}\}_{i=1}^{N}$ with a corresponding language description $l^{k}$ guiding the goal of the robot. Each trajectory $\tau^{k}_{i}$ in expert demonstrations with length $L^{k}<H$ consists of state-action pairs $\{(s_{t}^{k},a_{t}^{k})\}_{t=1}^{L^{k}}$ sampled from policy $\pi$ conditioned with task $T^{k}$. The policy is trained with a behavioral cloning loss [237] that aims to mimic the action demonstrations in the dataset:
@@
Benchmarks. To evaluate the performance of robots in continual learning, many benchmarks have been constructed in continual learning scenarios from reinforcement learning to vision language action [238, 239, 124]. For example, Continual World [238] builds a continual reinforcement learning benchmark on robotic manipulation tasks. It provides a realistic, computationally accessible testbed that highlights the limitations of existing continual learning methods in reinforcement learning (RL) settings, which often struggle with forward transfer despite mitigating forgetting. LIBERO [124] builds an imitation learning environment for robots to analyze the continual learning problem of robots from many factors, such as algorithms, architectures, task sequences, variations, and so on, which provides a comprehensive environment for subsequent work.
@@ 5.2 Architecture-based Approach
Architecture-based approaches investigate continual learning through structural design, with a focus on enhancing skill representation and reuse. One line of work explores skill-aware architectures that enable models to distinguish, merge, and update skills [240, 126, 125, 127], or construct extensible modular designs for skill composition [241, 242]. LOTUS [125] proposes a hierarchical imitation learning framework that incrementally builds a skill library via unsupervised segmentation of human demonstrations using open-vocabulary vision models. It updates or adds new skills through clustering and composes them via a meta-controller, achieving superior performance in lifelong manipulation through continual discovery and experience replay. IsCiL [243] introduces a prototype-based memory system that incrementally learns shareable skills for adaptation in dynamic environments. By mapping demonstrations to state embeddings and retrieving relevant skill prototypes, it employs adapter modules to preserve prior knowledge. LEGION [127] presents a lifelong reinforcement learning framework that combines Bayesian non-parametric clustering with language-conditioned embeddings. Using a Dirichlet Process Mixture Model (DPMM), it dynamically organizes skills and semantically aligns task understanding, enabling scalable skill accumulation without performance degradation. RLCM [244] formulates a modular lifelong RL framework that composes policies from learned neural modules. It adopts a two-phase learning process—online module selection and adaptation for new tasks, and offline consolidation via batch RL and experience replay—to retain and generalize skills over time.
Another line of research focuses on enhancing robotic action through architectural designs that leverage Large Language Models (LLMs) for high-level guidance in tasks such as planning [245] and skill demonstration [246, 247]. These approaches utilize pre-trained LLMs as knowledge bases, retrieving relevant information via textual prompts to facilitate skill acquisition and enable the transfer of prior knowledge to new tasks. Owing to their data efficiency and rapid inference capabilities, LLMs demonstrate strong potential across diverse robotic scenarios.
@@ 5.3 Regularization-based Approach
Regularization methods for continual learning usually try to keep important parameters or features unchanged and impose feature-level or parameter-level regularization during training, therefore maintaining the ability for past skills [248, 128, 129, 243]. CAMA [128] introduces Confidence-Aware Moving Average, a method that dynamically updates stored knowledge (logits) using confidence scores to balance past and current information without requiring task boundaries. PolyTask [248] proposes a method combining task-specific policy training with behavior distillation, which consolidates individual policies into a single model using offline data. With demonstration-guided reinforcement learning and distilling policy outputs, it handles continual actions and lifelong learning by avoiding interference between tasks. M2Distill [129] proposes a multi-modal distillation framework that preserves consistency across vision, language, and action modalities between incremental learning steps. It exploits: (1) multi-modal feature distillation to constrain latent space drift; and (2) policy distillation to maintain alignment of Gaussian mixture model (GMM) action distributions. The approach ensures skill retention without storing prior data, advancing lifelong learning for real-world robotic applications. ELIRL [249] introduces Lifelong Inverse Reinforcement Learning (Lifelong IRL), addressing the challenge of learning multiple sequential tasks and proposing Efficient Lifelong IRL to enable knowledge transfer across tasks, improving performance while minimizing user-provided data.
@@ 5.4 Replay-based Approach
Due to the spatial and temporal complexity of robotic manipulation in various and dynamic scenarios, it has become common practice to store part of previous samples and replay them [250, 38] during training procedure to retain the knowledge of previous tasks [130, 127, 241, 131]. iManip [130] addresses the challenge of catastrophic forgetting in robotic manipulation. Given that traditional methods fail to retain previously learned skills, ignoring temporal complexity and action complexity, the introduced framework integrates a temporal replay strategy to preserve task continuity and an extendable PerceiverIO [251] architecture with adaptable action prompts. Experimental results highlight the efficiency, scalability, and robustness across complex, long-horizon tasks. Decision-RWKV [241] exploits a novel sequence modeling framework that integrates the efficiency of the Receptance Weighted Key Value (RWKV) with experience replay mechanisms. It leverages the linear complexity and memory retention capabilities of RWKV to enable efficient lifelong learning across diverse manipulation tasks, and the experience replay combines a replay buffer for historical experiences reuse while acquiring new skills. RWLA [131] combines retrieval-based local adaptation with a selective weighting mechanism to restore forgotten skills. The method retrieves relevant past demonstrations from a memory buffer, identifies critical failure-prone segments, and prioritizes these segments during local fine-tuning, which allows robots to adapt continually while maintaining performance across diverse manipulation tasks.
@@ 6.1 Problem Setup
where $z_{t}$ denotes the latent code of an image with the new concept, and $z_{t}^{p}$ corresponds to a class-specific image used for prior preservation. $c$ and $c^{p}$ are the respective conditioning vectors; for instance, $c$ may be “a photo of [$\mathcal{V}^{*}$] dog”, while $c^{p}$ is “a photo of dog”. The special token [$\mathcal{V}^{*}$] is a unique identifier representing the user-specific concept learned from the provided images. $\lambda$ controls the regularization strength.
Continual learning in diffusion models aims to go beyond single-concept personalization by enabling the incremental acquisition of new concepts while avoiding catastrophic forgetting and eliminating the need to retrain on previous data. Specifically, as shown in Figure 7, assume a user intends to continuously generate content based on newly introduced specific concepts. The generation tasks are defined as $\mathcal{T}={\{\mathcal{T}^{k}}\}_{k=1}^{K}$ with corresponding datasets ${\{\mathcal{D}^{k}}\}_{k=1}^{K}$. Each $k$-th generation task consists of $n_{k}$ pairs of concept images $x_{i}^{k}$ and prompts $c_{i}^{k}$, where $K$ denotes the total number of tasks, and $n_{k}$ is typically in the range of 3 to 5 in accordance with existing studies. Accordingly, to learn the $k$-th personalized concept, the model is optimized using the following training loss:
@@ 6.3 Regularization-based Approach
MuseumMaker [133] aims to retain the knowledge of past learned styles while incorporating new customized styles. To mitigate catastrophic overfitting and maintain the integrity of original concepts in diffusion models, it introduces a style distillation loss that captures disentangled style representations. Furthermore, it proposes a dual regularization scheme for LoRA training, which constrains model optimization from both the weight and feature perspectives to achieve balanced knowledge integration. LFS-Diffusion [134] proposes a data-free knowledge distillation strategy applied across diffusion time steps during training, aiming to constrain the model from overfitting to current personalized concepts and thereby preserve the original generative capabilities. During inference, an in-context generation mechanism is introduced to mitigate forgetting of previously learned knowledge and ensure better retention of past concepts. CIDM [135] introduces a concept consolidation loss that employs learnable layer-wise concept tokens and an orthogonal subspace regularizer to disentangle task-specific and task-shared knowledge. During inference, an elastic weight aggregation module merges low-rank weights of previously learned concepts based on their importance. To alleviate concept neglect in multi-concept composition, CIDM further proposes a context-controllable synthesis strategy that enhances regional feature expressiveness via layer-wise textual embeddings and guides generation through region-aware noise estimation. ConceptGuard [136] introduces a set of lightweight yet effective strategies to mitigate forgetting in continual concept learning for diffusion models. It employs shift embeddings for dynamic concept updates, concept-binding prompts to disentangle and compose concepts over time, and memory preservation regularization to slow down harmful parameter drift. Additionally, a priority queue adaptively manages concept replay based on their importance and temporal order. Jha et al. [137] utilize diffusion classifier scores as implicit class-conditional density estimators. This method simultaneously regularizes both parameter-space and function-space consolidation, enabling effective adaptation to evolving user concepts.
@@ 7 Future Directions
Learning Paradigm for Continual Learning. While conventional continual learning methods predominantly rely on supervised fine-tuning (SFT) for incremental knowledge acquisition, emerging reinforcement learning (RL)-based training paradigms, particularly exemplified by GRPO [270], have demonstrated stronger generalization capabilities across a variety of model architectures. This performance advantage has been consistently observed not only in large language models [271, 272, 273] but also in multimodal large language models [274, 275, 276, 277] and vision-language action models [278, 279], suggesting that RL-based methods may serve as a more robust alternative to traditional SFT strategies in continual learning settings. Recent studies [280] have revealed a fundamental difference between the two paradigms: SFT tends to overfit specific training distributions through pattern memorization, whereas RL emphasizes structural comprehension of task objectives, enabling better generalization to novel scenarios. This distinction provides strong motivation for exploring continual learning within an RL framework, which may naturally support long-term adaptation, task abstraction, and reinforcement-driven retention strategies.
@@
Self-generation for Continual Learning. A core challenge in continual learning lies in the inaccessibility of previous task data. In discriminative models, a widely adopted solution involves training auxiliary generative models (e.g.,, GANs [281] or diffusion models) to synthesize pseudo-samples of prior tasks, thereby mitigating catastrophic forgetting [282, 283, 284, 285]. While effective, this approach introduces significant computational overhead and may suffer from a distributional mismatch between generated and original data. In contrast, generative models offer a more integrated solution by leveraging their inherent capacity to reconstruct and regenerate data. These models can autonomously replay past task distributions during the learning of new tasks, without the need for separate generative modules. This capability offers two major benefits: reducing additional architectural complexity and achieving better alignment with the original data distribution. As a result, self-generation in generative models presents a promising direction for building more efficient, scalable, and distributionally consistent continual learning systems.
@@
Continual Learning in Large Scale Models. Most existing studies on continual learning for generative models have focused on small- to medium-scale architectures (e.g., models with up to 7B and 13B parameters). However, emerging evidence [286, 287] suggests that continual learning behaviors can differ substantially as the model scale increases. This observation highlights two critical research directions. First, there is a pressing need to establish specialized evaluation benchmarks that mitigate information leakage risks [288], particularly given the large and often proprietary datasets used to train large-scale models. Second, it is essential to explore parameter-efficient knowledge integration strategies—such as Chain-of-Thought reasoning (CoT)[166, 289], In-Context Learning (ICL)[290, 291, 292], Agents [293, 294, 295], and Retrieval-Augmented Generation (RAG) [296, 297]—as conventional fine-tuning becomes increasingly infeasible for models with extensive parameter counts.
@@
Continual Learning for Multimodal Generation Beyond Vision and Language. While most existing research in continual generative learning focuses on vision-language tasks, real-world multimodal systems are increasingly incorporating richer modalities such as audio [298], 3D scenes [299], and motion data [300]. In such settings, modality-continual learning becomes a critical yet underexplored challenge. For example, a pre-trained MLLM may be required to incorporate speech input or generate spoken responses. Without appropriate continual learning mechanisms, learning the new audio modality could result in severe forgetting of previously acquired vision-language capabilities. Similarly, in VLA models deployed in robotics or embodied agents, the integration of new sensor modalities (e.g., depth and haptic signals) necessitates stable expansion of generative capacity without overwriting existing knowledge. These scenarios highlight the need for continual learning methods capable of integrating new modalities while preserving alignment and generative quality across previously learned ones.